### v4

@@ Appendix A  Derivation of the Design-Based Likelihood
$={\mathbb{P}}\Big(I_{11}+I_{10}=x_{I1},\ I_{11}+I_{01}=x_{C1}-\theta_{11}-\theta_{01},$
$\qquad\qquad I_{00}+I_{01}=x_{I0},\ I_{00}+I_{10}=x_{C0}-\theta_{00}-\theta_{10}\mid{\boldsymbol{\theta}}\Big).$
A realization of the data $\boldsymbol{X}$} may be produced from more than one realization of the distribution of potential outcomes in intervention $\boldsymbol{I}$}. To find the probability of a realization of $\boldsymbol{X}$}, we sum together the probabilities of each realization of $\boldsymbol{I}$} that could have produced it. We can index these realizations through the realization $i$ of $I_{11}$ and solve the following system of equations for the elements of $\boldsymbol{I}$}:
@@
$I_{11}$ $=i$
$I_{10}$ $=x_{I1}-i$
$I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1}-i$
$I_{00}$ $=x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}.$
The value $i$ is restricted to the set $\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})$ such that $\boldsymbol{I}$} remains within the support implied by $\boldsymbol{\theta}$}, namely $0\leq I_{11}\leq\theta_{11}$, $0\leq I_{10}\leq\theta_{10}$, $0\leq I_{01}\leq\theta_{01}$, and $0\leq I_{00}\leq\theta_{00}$. The probability of a realization of $\boldsymbol{X}$} is just the sum of the probability of each of these realizations of $\boldsymbol{I}$}:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})={\mathbb{P}}\big({\boldsymbol{X}}={\boldsymbol{x}}\mid{\boldsymbol{\theta}}\big)=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}{\mathbb{P}}\Big(I_{11}$ $=i,$
$I_{10}$ $=x_{I1}-i,$
$I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1}-i,$
$I_{00}$ $=x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}\mid{\boldsymbol{\theta}}\Big).$ (7)
@@
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}}$
@@
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$. (9)
where we have substituted $m=x_{I1}+x_{I0}$. In either case, once conditioning on the realized data $\boldsymbol{x}$}, the likelihood is proportional to
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $\propto\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$. (10)
This final expression is equivalent to the likelihood in copas1973.
Appendix B  Derivation of a Canonical Sampling-Based Likelihood
Suppose that subjects are sampled into the experiment with replacement from a population: let $q_{11}$ represent the share of the population who are always takers, $q_{10}$ the share of compliers, $q_{01}$ the share of defiers, and $q_{00}$ the share never takers. Sampling with replacement implies that each subject’s pair of potential outcomes is independently and identically distributed (IID) according to these shares.
@@ Appendix B  Derivation of a Canonical Sampling-Based Likelihood
This likelihood is equivalent to a likelihood appearing in barnard1947.
@@
This likelihood is equivalent to likelihoods from barnard1947 and kline2020.
@@
Appendix C  Bayes Optimality of Our Proposed Rule
Appendix C.1  Preliminaries on Statistical Decision Theory
Suppose a decision maker wishes to guess the joint distribution of potential outcomes in the sample. We write the decision maker’s guess as $\widehat{{\boldsymbol{\theta}}}$, and we define a utility function over a guess $\widehat{{\boldsymbol{\theta}}}$ and the true distribution $\boldsymbol{\theta}$}. We focus on the following utility specification, which yields one util when the guess is correct and zero utils when the guess is incorrect:
$u(\widehat{{\boldsymbol{\theta}}},{\boldsymbol{\theta}})=\mathbf{1}_{\{\widehat{{\boldsymbol{\theta}}}={\boldsymbol{\theta}}\}}.$
We also allow the decision maker to choose a randomized guess, which ascribes a probability distribution over the possible values of $\boldsymbol{\theta}$}. We define the decision maker’s utility over a randomized guess $\rho$ as the expected utility of guessing according to the probabilities ascribed by $\rho$:
@@ Appendix C.1  Preliminaries on Statistical Decision Theory
The decision maker chooses a decision rule that maps the observable data into (possibly) randomized guesses. [^3] We conflate here the standard definitions of “randomized decision rules” and “behavioral decision rules” (ferguson1967) for expositional clarity. In settings of perfect recall, such as the setting we study here, the space of randomized and behavioral decision rules is equivalent (kuhn1953). We write such a rule as $f:\boldsymbol{\mathcal{X}}\to\Delta({\boldsymbol{\Theta}})$, where $\boldsymbol{\mathcal{X}}$ is the space of possible data realizations, $\boldsymbol{\Theta}$} is the space of possible distributions of potential outcomes, and $\Delta({\boldsymbol{\Theta}})$ is the space of distributions over $\boldsymbol{\Theta}$}. Given a true distribution of potential outcomes $\boldsymbol{\theta}$}, the decision maker’s expected utility from following a decision rule $f$ is the expected value of $U(f({\boldsymbol{X}}),{\boldsymbol{\theta}})$ with respect to the experimental outcome $\boldsymbol{X}$}:
$EU(f,{\boldsymbol{\theta}})$ $=\mathbb{E}\big[U\big(f({\boldsymbol{X}}),{\boldsymbol{\theta}}\big)\mid{\boldsymbol{\theta}}\big]$
@@
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})f({\boldsymbol{x}})({\boldsymbol{\theta}}).$
The last line follows from our chosen utility function, under which the decision maker’s expected utility is equal to the probability of guessing correctly.
One sense in which a decision rule can be optimal is in terms of Bayes expected utility: the average expected utility obtained according to some subjective prior distribution over $\boldsymbol{\theta}$}:
@@
where $\pi\in\Delta({\boldsymbol{\Theta}})$ is a subjective prior belief about $\boldsymbol{\theta}$}. Decision rules that maximize this criterion are said to be “Bayes optimal.”
Appendix C.2  Conditions for Bayes Optimality

### v5

@@ Appendix A.1  Design-Based Likelihoods Can Vary Conditional on the Marginal Distributions of Potential Outcomes
$={\mathbb{P}}\Big(I_{11}+I_{10}=x_{I1},\ I_{11}+I_{01}=\theta_{11}+\theta_{01}-x_{C1},$
$\qquad\qquad I_{00}+I_{01}=x_{I0},\ I_{00}+I_{10}=\theta_{00}+\theta_{10}-x_{C0}\mid{\boldsymbol{\theta}}\Big).$
A realization of the data $\boldsymbol{x}$} may be produced from more than one realization of the joint distribution of potential outcomes in intervention $\boldsymbol{i}$}. To find the probability of the realization $\boldsymbol{x}$}, we sum together the probabilities of each $\boldsymbol{i}$} that could have produced it. Using $j$ as the realized number of always takers randomized into intervention $i_{11}$, we solve the following system of equations for the elements of the joint distribution of potential outcomes in intervention $\boldsymbol{I}$}:
@@
$I_{11}$ $=j$
$I_{10}$ $=x_{I1}-j$
$I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1}-j$
$I_{00}$ $=x_{I0}+x_{C1}+j-\theta_{11}-\theta_{01}.$
The value $j$ is restricted to the set $\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})$ such that $\boldsymbol{I}$} remains within the support implied by $\boldsymbol{\theta}$}, namely $0\leq I_{11}\leq\theta_{11}$, $0\leq I_{10}\leq\theta_{10}$, $0\leq I_{01}\leq\theta_{01}$, and $0\leq I_{00}\leq\theta_{00}$. The probability of a realization of $\boldsymbol{X}$} is just the sum of the probability of each of these realizations of $\boldsymbol{I}$}:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})={\mathbb{P}}\big({\boldsymbol{X}}={\boldsymbol{x}}\mid{\boldsymbol{\theta}}\big)=\sum_{j\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}{\mathbb{P}}\Big(I_{11}$ $=j,$
$I_{10}$ $=x_{I1}-j,$
$I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1}-j,$
$I_{00}$ $=x_{I0}+x_{C1}+j-\theta_{11}-\theta_{01}\mid{\boldsymbol{\theta}}\Big).$ (7)
@@
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{j\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{j}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-j}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-j}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+j-\theta_{11}-\theta_{01}}$
@@
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{j\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{j}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-j}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-j}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+j-\theta_{11}-\theta_{01}}\bigg/\binom{n}{m}$. (9)
In either case, once conditioning on the realized data $\boldsymbol{x}$}, the likelihood is proportional to
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $\propto\sum_{j\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{j}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-j}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-j}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+j-\theta_{11}-\theta_{01}}$. (10)
This final expression appears in (2) and is equivalent to the likelihood in Copas (1973). This expression also represents the numerator of the likelihood values we calculate in the stylized example in Section 1.1, which vary conditional on the marginal distributions of potential outcomes.
Appendix A.2  Canonical Sampling-Based Likelihoods are Flat Conditional on the Marginal Distributions of Potential Outcomes
Suppose that subjects are sampled into the experiment with replacement from a population. Let $q_{11}$ represent the share of the population who are always takers, $q_{10}$ the share of compliers, $q_{01}$ the share of defiers, and $q_{00}$ the share of never takers. Sampling with replacement implies that each subject’s pair of potential outcomes is independently and identically distributed (IID) according to these shares.
@@
This likelihood is equivalent to a likelihood appearing in Barnard (1947).
@@
This likelihood is equivalent to likelihoods from Barnard (1947) and Kline and Walters (2021).
@@
Appendix A.3  Alternative Sampling-Based Likelihoods Can Vary Conditional on the Marginal Distributions of Potential Outcomes
When subjects are sampled from a finite population without replacement, their potential outcomes are no longer independently and identically distributed, and the sampling-based likelihood maintains some of the curvature found in the design-based likelihood. Let the finite population consist of ${\boldsymbol{s}}\equiv(s_{11},s_{10},s_{01},s_{00})$ where $s_{11}$ denotes the number of always takers, $s_{10}$ the number of compliers, $s_{01}$ the number of defiers, and $s_{00}$ the number of never takers. Let the total population size be $k=s_{11}+s_{10}+s_{01}+s_{00}$. Suppose that $n$ subjects are sampled without replacement from this population. As before, let ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ denote the counts of each type in the experimental sample, where $\theta_{11}+\theta_{10}+\theta_{01}+\theta_{00}=n$. Under sampling without replacement, the distribution of $\boldsymbol{\theta}$} conditional on the population counts $\boldsymbol{s}$} follows the multivariate hypergeometric distribution:
$\mathbb{P}({\boldsymbol{\theta}}\mid{\boldsymbol{s}})=\frac{\binom{s_{11}}{\theta_{11}}\binom{s_{10}}{\theta_{10}}\binom{s_{01}}{\theta_{01}}\binom{s_{00}}{\theta_{00}}}{\binom{k}{n}}$.
Among the sampled subjects, the experiment randomly assigns subjects to intervention and control. Because the experimental sample itself is random, the likelihood of the population counts $\boldsymbol{s}$} must integrate over all possible realizations of the sample type counts $\boldsymbol{\theta}$}. Applying the law of total probability yields
$\mathcal{L}({\boldsymbol{s}}\mid{\boldsymbol{x}})=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})\mathbb{P}({\boldsymbol{\theta}}\mid{\boldsymbol{s}}).$
Note that, conditional on the sample distribution $\boldsymbol{\theta}$}, the experimental outcome $\boldsymbol{x}$} does not depend on the entire population distribution $\boldsymbol{s}$}. Substituting the expression for $\mathbb{P}({\boldsymbol{\theta}}\mid{\boldsymbol{s}})$ from above gives
$\mathcal{L}({\boldsymbol{s}}\mid{\boldsymbol{x}})=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})\frac{\binom{s_{11}}{\theta_{11}}\binom{s_{10}}{\theta_{10}}\binom{s_{01}}{\theta_{01}}\binom{s_{00}}{\theta_{00}}}{\binom{k}{n}}$,
where $\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ is the design-based likelihood of the experiment sample, which is given by equation (8) under simple randomization or equation (9) under complete randomization, both of which can be found in Appendix A.1.
Unlike the population likelihoods under sampling with replacement we derive in Appendix A.2, which depend on the joint distribution of potential outcomes only through the marginal distributions, the population likelihood under sampling without replacement is a function of the entire joint distribution of potential outcomes. As a result, the population likelihood under sampling without replacement varies within the population Fréchet bounds on defiers: if 6 people are drawn from a finite population of 24 without replacement, 2 of 3 randomized into intervention take up, and 1 of 3 randomized into control take up, then the finite population consisting of 8 always takers, 8 compliers, and 8 never takers has a likelihood of 0.0374, while the finite population consisting of 16 compliers and 8 defiers has a likelihood of 0.0548, despite the fact that these joint distributions have the same marginal distributions.
Appendix B  Justifications for Our Maximum Likelihood Decision Rule
Appendix B.1  Preliminaries on Statistical Decision Theory
Maximum likelihood estimates are ubiquitous in statistics and econometrics, and we offer three justifications for their use in our design-based model of an experiment. The first two justifications are common in statistical decision theory, which treats estimation and inference as decision problems under uncertainty and looks for optimal estimators—or “decision rules.” Suppose a decision maker wishes to guess the joint distribution of potential outcomes in the sample. We write the decision maker’s guess as $\widehat{{\boldsymbol{\theta}}}$, and we define a utility function $u$ over a guess $\widehat{{\boldsymbol{\theta}}}$ and the true distribution $\boldsymbol{\theta}$}. We also allow the decision maker to choose a randomized guess, which ascribes a probability distribution over the possible values of $\boldsymbol{\theta}$}. We define the decision maker’s utility over a randomized guess $\rho$ as the expected utility of guessing according to the probabilities ascribed by $\rho$:
@@ Appendix B.1  Preliminaries on Statistical Decision Theory
The decision maker chooses a decision rule that maps the observable data into (possibly) randomized guesses. [^1] We conflate here the standard definitions of “randomized decision rules” and “behavioral decision rules” (Ferguson, 1967) for expositional clarity. In settings of perfect recall, such as the setting we study here, the space of randomized and behavioral decision rules is equivalent (Kuhn, 1953). We write such a rule as $f:\boldsymbol{\mathcal{X}}\to\Delta$, where $\boldsymbol{\mathcal{X}}$ is the space of possible data realizations, $\boldsymbol{\Theta}$} is the space of possible distributions of potential outcomes, and $\Delta$ is the space of distributions over $\boldsymbol{\Theta}$}. Given a true distribution of potential outcomes $\boldsymbol{\theta}$}, the decision maker’s expected utility from following a decision rule $f$ is the expected value of $U(f({\boldsymbol{X}}),{\boldsymbol{\theta}})$ with respect to the experimental outcome $\boldsymbol{X}$}:
@@
One sense in which a decision rule can be optimal is in terms of Bayes expected utility, which is the average expected utility obtained according to some subjective prior distribution over $\boldsymbol{\theta}$}:
@@
where $\pi\in\Delta$ is a subjective prior belief about $\boldsymbol{\theta}$}. Decision rules that maximize this criterion are said to be “Bayes optimal.”
To define the maximum likelihood decision rule, let $\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})$ be the set of joint distributions $\boldsymbol{\theta}$} that maximize the likelihood given the observed data $\boldsymbol{x}$}:
where $\{\bullet\}$ is the counting measure. When the likelihood is unimodal, the decision rule chooses the maximizer with probability one.
Appendix B.2  Our Maximum Likelihood Decision Rule Varies Systematically with the Data
Our first justification for our maximum likelihood decision rule is that it varies with the data, so it is “reasonable” as argued by Ferguson (1967) in his introduction to statistical decision theory. Direct inspection of the likelihood in (2) does not reveal obvious patterns in the maximum likelihood estimate. However, by computing the maximum likelihood estimate for every possible realization of the data in fixed sample sizes, we can observe clear patterns in the estimates. Figure B.1 illustrates how the maximum likelihood estimates from our proposed decision rule vary systematically with all possible realizations of the data in completely randomized experiments of 50 and 200 subjects with half in intervention. In a sample of 200 with half in intervention, there are $10,201=101\times 101$ possible realizations of the data and $1,373,301={203\choose 3}$ possible joint distributions of potential outcomes, so the visualization reflects results of grid searches over $14,009,043,501=10,201\times 1,373,301$ total elements.
Figure B.1: Visualization of Our Proposed Statistical Decision Rule: How the MLE Varies with All Possible Data from Samples of 50 and 200 with Half in Intervention
In the convex lens-shaped region around the upward-sloping diagonal, MLEs only include always and never takers: the two types in the null hypothesis of the Fisher (1935) exact test. In samples of 50 and 200, the region in which Fisher’s exact test fails to reject the null hypothesis at the 5% level is depicted within the grey dashed line. It is larger than the region in which the MLE includes only always and never takers in both sample sizes.
The MLEs include three types in all other regions, with two regions for each combination. Even though the maximizer of the likelihood only includes two types in our stylized example from Figure 1, maximization of the likelihood sometimes requires trading off between the higher likelihood of fewer types and the higher likelihood of more balance between intervention and control within each type. If restricting the sample to two types requires too much imbalance between intervention and control, the maximizer of the likelihood includes three types. [^2] Viewing Pascal’s triangle as a representation of $N$ choose $K$ can help us to visualize the tradeoff between fewer types (moving to the next row of the triangle, increasing $N$ within each type) and more balance within each type (moving to the middle of the triangle within a row, moving $K$ closer to $N/2$ within each type). We thank Elizabeth Ananat for this point. There are no regions in which a distribution that maximizes the likelihood contains all four types. Therefore, many joint distributions of potential outcomes never maximize the likelihood for any realization of the data. For example, if a sample of 50 consisted of 18 always takers, 22 compliers, 8 defiers, and 2 never takers, the most probable realization of the data would be takeup of 20 in intervention and 13 in control, but the corresponding MLE estimate of 26 always takers, 14 compliers, 0 defiers, and 10 never takers would miss the 8 defiers entirely. We have observed MLEs that contain all four types in experiments with intervention and control arms unequal in size, although we have only observed them tied with MLEs that contain three types.
The region outlined in bold, in which the estimated average effect is positive, is of particular interest because MLEs in the shaded portion of the region include both defiers and compliers, contrary to monotonicity assumptions, as in the stylized example in Figure 1. In all even-sized samples up to 200, the MLE includes defiers when takeup is below half in control and above half in intervention, unless takeup is zero in control or full in intervention. The experiment depicted in Figure 1 falls into the region with defiers because the takeup rate is 1/3 in control and 2/3 in intervention. As shown, the MLE always includes defiers when the estimated average effect is greater than half, unless takeup is zero in control or full in intervention, and it can include defiers even when Fisher’s exact test rejects at the 5% level. This pattern might seem surprising or even disconcerting because at the edges of the shaded regions, small changes in the data can yield large changes in the estimated numbers of defiers. However, while point estimates can change abruptly, credible sets on either side of a transition overlap substantially, reflecting the underlying continuity of the likelihood. We recommend that applied researchers report credible sets alongside maximum likelihood estimates, as we do in our applications. Moreover, the persistence of the pattern in samples of 50 and 200 suggests that the region in which the MLE includes defiers will persist in larger samples.
The shading demonstrates that the MLE reveals evidence beyond the average effect: holding the estimated average effect constant along upward sloping diagonal lines, MLEs can include zero or many defiers. Copas (1973) claims that the likelihood is always maximized at a joint distribution of potential outcomes that preserves the direct estimates of the marginal distributions of potential outcomes from equation (4) and which has either the maximal or minimal number of defiers. In practice, we often find that the likelihood is maximized at a distribution that deviates slightly from these marginal estimates.
Appendix B.3  Our Maximum Likelihood Decision Rule is Bayes Optimal
We Motivate Conditions for Bayes Optimality:
Our second justification for our proposed maximum likelihood decision rule is that it is Bayes optimal under a specific utility function and a specific prior. Bayes optimal decision rules can be derived for alternative utility functions and priors through the numerical solution of a linear programming problem, but we offer motivations for the specifications we have chosen, acknowledging their subjectivity. For the utility function, we specify a discrete utility function that assigns zero utility for incorrect guesses and an arbitrary amount of strictly positive utility for correct guesses:
$u(\widehat{{\boldsymbol{\theta}}},{\boldsymbol{\theta}})=\mathbf{1}_{\{\widehat{{\boldsymbol{\theta}}}={\boldsymbol{\theta}}\}}.$ (15)
For the prior, we specify a prior that is uniform over all possible distributions of potential outcomes for $n$ subjects:
$\pi({\boldsymbol{\theta}})=\frac{1}{\#\{{\boldsymbol{\Theta}}\}},\ \forall{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}.$ (16)
One motivation for our utility function is that we can interpret it as the payoff function in a game that rewards “being correct,” as in the well-known Monty Hall game, where the contestant receives a prize only if they guess the correct door. As we review below, the maximum likelihood decision rule under this utility function is the maximum a posteriori (MAP) rule, which chooses the mode of the posterior distribution. Under a uniform prior, the posterior distribution is proportional to the likelihood, and the MAP estimate equals the MLE estimate.
The desirable feature of our uniform prior is that it assigns equal weight to all distributions in the same Fréchet set, so it is “uninformative” in the sense that it does not confer any information about the joint distribution of potential outcomes beyond their marginal distributions. This uninformative prior is consistent with the beliefs of a statistician who places equal weight on all values of defiers within the estimated Fréchet bounds. Furthermore, we can view our uniform prior as the result of sampling from an underlying population over which we hold an uninformative hyperprior. In sampling settings, it is common to assume a Dirichlet hyperprior over population shares. We choose a Dirichlet hyperprior over the population shares of always takers, compliers, defiers, and never takers with all concentration parameters set to one, which is equivalent to an uninformative uniform distribution for the population. Under standard assumptions on sampling from the population, the potential outcomes of subjects are IID, and the sample distribution of potential outcomes follows a multinomial law conditional on the population distribution. Combining the Dirichlet hyperprior and the sampling process yields a Dirichlet-multinomial prior over the counts of each type in the sample. The Dirichlet-multinomial prior with concentration parameters set to one is equivalent to our uniform prior. In our stylized example in Figure 1, our uniform prior implies that the prior is the same in all three rows, so the updated posterior is proportional to the likelihood.
Even for other prior specifications, the ratio of the beliefs between any two distributions updates as long as the likelihood ratio differs from one, including within Fréchet sets. For example, consider an alternative prior that is uniform over each type for each person in the sample. This prior arises from a highly informative degenerate hyperprior that assumes the population shares of each type are equal. The resulting prior over sample distributions varies within the sample Fréchet sets, and is therefore also informative. In our stylized example in Figure 1, the resulting prior is 0.0220 for the first row, 0.0293 for the second row, and 0.0037 for the third row. The updated posterior distribution maintains curvature within the Fréchet set and the ratio of the prior to the posterior is highest for the third row that contains two defiers. To achieve a flat posterior distribution across the rows depicted in Figure 1, the prior would have to be very specific and informative: the prior for the first row with no defiers would have to be exactly 0.75 (=(20/8)/(20/6)) times as large as the second row with one defier and 1.5 (=(20/8)/(20/12)) times as large as the prior for the third row with two defiers.
A Standard Proof Establishes Bayes Optimality:
Proposition 1 establishes the Bayes optimality of the maximum likelihood decision rule under the conditions described above. [^3] We thank Andriy Norets and Thomas Wiemann for bringing the standard conditions that establish Bayes optimality to our attention.
Proposition 1.
When utility $u$ takes the form in (15) and the prior belief $\pi$ takes the form in (16), the maximum likelihood decision rule $\widehat{{\boldsymbol{\theta}}}_{MLE}$ is Bayes optimal:
$\widehat{{\boldsymbol{\theta}}}_{MLE}\in{\arg\max}_{f}EU_{\pi}(f).$
Proof.